Two packages together weigh 32 kilograms the smaller package weighs 6 kilograms less than the larger. How much does each package weigh? Use the variable s and L show setup and work. This is what I have so far but I am confused now...
s + L = 32
-2(s+L=32)
-2s+-2L= -2(32)
-2s -2L = -64
-2s-2L = -64
2s + 6L =

Answer by checkley77(7049)

s + L = 32
-2(s+L=32)
-2s+-2L= -2(32)
-2s -2L = -64
-2s-2L = -64
2s + 6L =
My solution would be:
S+L=32 OR S=32-L
S=L-6
NOW SET THESE 2 EQUATIONS EQUAL TO EACH OTHER.
S=S OR:
L-6=32-L
L+L=32+6
2L=38
L=38/2
L=19 KG. FOR THE LARGER PACKAGE.
S+19=32
S=32-19
S=13 KG. FOR THE SMALLER PACKAGE.